# Fine-grained LLM Agent: Pinpointing and Refining Large Language Models via Fine-Grained Actionable Feedback

### v4

@@ Abstract
Recent large language models (LLM) are leveraging human feedback to improve their generation quality. However, human feedback is costly to obtain, especially during inference. In this work, we propose Fine-grained LLM Agent, an inference time optimization method to refine LLM’s output. The core idea is to use a learned fine-grained feedback model to pinpoint defects and guide LLM to refine them iteratively. Using original LLM as a proposal of edits, Fine-grained LLM Agent searches for defect-less text via simulated annealing, trading off the exploration and exploitation. We conduct experiments on three text generation tasks, including machine translation, long-form question answering (QA), and topical summarization. Fine-grained LLM Agent consistently outperforms all baseline approaches, achieving improvements up to 1.7 MetricX points on translation tasks, 8.1 ROUGE-L on ASQA, 2.2 ROUGE-L on topical summarization.
@@ 1 Introduction
Figure 1: An overview of our Fine-grained LLM Agent: We start from LLM’s initial generation and iteratively refine the generation, based on fine-grained actionable feedback. We use a simulated annealing technique to accept or reject the proposed revision at each step.
@@
In this work, we propose Fine-grained LLM Agent, an inference-time optimization method to improve the quality of generated text. Our Fine-grained LLM Agent starts with LLM’s initial output, then uses a learned error pinpoint model to provide fine-grained feedback about the location and type of defects in the text. We then use a refinement model (same or another LLM) to follow the feedback instruction and generate candidate text. The fine-grained feedback provides more much precise information about what exactly is wrong in the generated text, resulting in higher quality revision.
However, due to the large search space, the refinement model is imperfect; it often fails to correct all of the errors identified by the feedback model in one iteration Madaan et al. (2023). We formulate the iterative refinement procedure into a local search problem. It alternates between the feedback generation and refinement in multiple iterations, with the goal of searching for the highest scoring output according to the feedback model. To this end, we develop a simulated annealing technique in Fine-grained LLM Agent to trade off between exploring many possible edits and quickly reaching optimal text. Figure 1 shows overview of our approach.
We evaluate Fine-grained LLM Agent on three text generation tasks, including machine translation (WMT Kocmi et al. (2022)), long-form question answering (ASQA Stelmakh et al. (2022)) and topic summarization Saunders et al. (2022), because they have a large number of annotated outputs with fine-grained error spans Freitag et al. (2021a); Saunders et al. (2022); Wu et al. (2023). We use those ratings to train an automatic error pinpoint model that generates a list of error spans along with error categories and severities without the aid of a reference text (which is unavailable during inference) Fernandes et al. (2023); Xu et al. (2023b). This model serves as our substitute for human feedback. Our experiments show that Fine-grained LLM Agent results in higher-quality text compared to baseline methods using other feedback (scalar or binary score) or other search techniques. Our contributions are:
  • We propose Fine-grained LLM Agent, an inference time optimization method to iteratively refine LLM’s output with fine-grained actionable feedback, achieving best trade-off between search space and optimal quality.
  • We demonstrate that Fine-grained LLM Agent consistently outperforms all baseline approaches, achieving improvements up to 1.7 MetricX points on translationn tasks, 8.1 ROUGE-L on ASQA and 2.2 ROUGE-L improvements on topical summarization. Humans demonstrate a significant preference for the output of Fine-grained LLM Agent over the baseline outputs.
@@ 6 Results
Fine-grained LLM Agent 75.9 74.2 83.5 79.3 26.1 30.5
Table 3: We include three baseline models using coarse feedback: Improve, Binary-QE, Score-QE, BLEURT-Binary-QE, BLEURT-Score-QE and Fine-grained LLM Agent, which is guided by our fine-grained error pinpoint model. All results are obtained through greedy decoding. In Appendix Table 14 and 15, we report additional results on open sourced LLMs and results of COMET scores to demonstrate the effectiveness of our method on open sourced models.
@@ 6.1 Meta-Evaluating the Pinpoint Model
Model WMT’22 WMT’23 ASQA Summ Zh-En En-De Zh-En En-De QA Summ PaLM-2 0-shot 66.1 77.0 65.7 75.1 17.6 25.2 Improve 67.7 77.1 67.5 75.9 19.2 25.5 Score-QE 67.5 77.2 67.2 76.3 21.9 26.4 Binary-QE 67.6 77.9 67.3 76.5 21.6 26.0 Fine-grained LLM Agent 68.8 78.6 68.2 76.9 26.1 28.1
@@ 6.4 Iterative Refinement
Fine-grained LLM Agent vs Win Neutral Lose Win lose ratio 0-shot 38% 46% 16% 2.34 Improve 39% 45% 16% 2.44 BLEURT-SCORE-QE 41% 44% 15% 2.79 BLEURT-binary-QE 33% 48% 19% 1.76 score-qe 33% 40% 27% 1.23 binary-qe 34% 48% 18% 1.84
Table 5: We conduct head-to-head human evaluation on Fine-grained LLM Agent against all baselines for single step refinement at WMT22 En-De. We report percentages of win, neutral and lose and win/lose ratio in the table.
@@ 7 Conclusion
In this work, we proposed Fine-grained LLM Agent for incorporating a fine-grained error pinpoint into a text generation pipeline via refinement-feedback model loop. We empirically demonstrated that our fine-grained error pinpoint model with simulated annealing achieves superior performance compared to baseline feedback models and other search algorithms across three text generation tasks. Lastly, our experimental findings are further solidified by the human evaluation study. Humans demonstrate a significant preference for the output of Fine-grained LLM Agent over the baseline outputs.
@@ 8 Limitations
While Fine-grained LLM Agent in theory can be applied to landscape of instruction-fine-tuned large language models, it would be noteworthy to mention that the efficiency may differ when applied with models that lack instruction following capability. Training a large language model with instruction, feedback following ability can be great a future direction to mitigate this issue.
@@ Appendix B Additional Results
Table 14: Fine-grained LLM Agent’s performance on Mistral MoE Jiang et al. (2024) compared to coarse feedback, measured by metricX.
@@
Table 15: Fine-grained LLM Agent’s performance on PALM2 compared to coarse feedback, measured by COMET22.

### v5

# LLMRefine: Pinpointing and Refining Large Language Models via Fine-Grained Actionable Feedback

@@ Abstract
Recent large language models (LLM) are leveraging human feedback to improve their generation quality. However, human feedback is costly to obtain, especially during inference. In this work, we propose LLMRefine, an inference time optimization method to refine LLM’s output. The core idea is to use a learned fine-grained feedback model to pinpoint defects and guide LLM to refine them iteratively. Using original LLM as a proposal of edits, LLMRefine searches for defect-less text via simulated annealing, trading off the exploration and exploitation. We conduct experiments on three text generation tasks, including machine translation, long-form question answering (QA), and topical summarization. LLMRefine consistently outperforms all baseline approaches, achieving improvements up to 1.7 MetricX points on translation tasks, 8.1 ROUGE-L on ASQA, 2.2 ROUGE-L on topical summarization.
@@ 1 Introduction
Figure 1: An overview of our LLMRefine: We start from LLM’s initial generation and iteratively refine the generation, based on fine-grained actionable feedback. We use a simulated annealing technique to accept or reject the proposed revision at each step.
@@
In this work, we propose LLMRefine, an inference-time optimization method to improve the quality of generated text. Our LLMRefine starts with LLM’s initial output, then uses a learned error pinpoint model to provide fine-grained feedback about the location and type of defects in the text. We then use a refinement model (same or another LLM) to follow the feedback instruction and generate candidate text. The fine-grained feedback provides more much precise information about what exactly is wrong in the generated text, resulting in higher quality revision.
However, due to the large search space, the refinement model is imperfect; it often fails to correct all of the errors identified by the feedback model in one iteration Madaan et al. (2023). We formulate the iterative refinement procedure into a local search problem. It alternates between the feedback generation and refinement in multiple iterations, with the goal of searching for the highest scoring output according to the feedback model. To this end, we develop a simulated annealing technique in LLMRefine to trade off between exploring many possible edits and quickly reaching optimal text. Figure 1 shows overview of our approach.
We evaluate LLMRefine on three text generation tasks, including machine translation (WMT Kocmi et al. (2022)), long-form question answering (ASQA Stelmakh et al. (2022)) and topic summarization Saunders et al. (2022), because they have a large number of annotated outputs with fine-grained error spans Freitag et al. (2021a); Saunders et al. (2022); Wu et al. (2023). We use those ratings to train an automatic error pinpoint model that generates a list of error spans along with error categories and severities without the aid of a reference text (which is unavailable during inference) Fernandes et al. (2023); Xu et al. (2023b). This model serves as our substitute for human feedback. Our experiments show that LLMRefine results in higher-quality text compared to baseline methods using other feedback (scalar or binary score) or other search techniques. Our contributions are:
  • We propose LLMRefine, an inference time optimization method to iteratively refine LLM’s output with fine-grained actionable feedback, achieving best trade-off between search space and optimal quality.
  • We demonstrate that LLMRefine consistently outperforms all baseline approaches, achieving improvements up to 1.7 MetricX points on translationn tasks, 8.1 ROUGE-L on ASQA and 2.2 ROUGE-L improvements on topical summarization. Humans demonstrate a significant preference for the output of LLMRefine over the baseline outputs.
@@ 6 Results
LLMRefine 75.9 74.2 83.5 79.3 26.1 30.5
Table 3: We include three baseline models using coarse feedback: Improve, Binary-QE, Score-QE, BLEURT-Binary-QE, BLEURT-Score-QE and LLMRefine, which is guided by our fine-grained error pinpoint model. All results are obtained through greedy decoding. In Appendix Table 14 and 15, we report additional results on open sourced LLMs and results of COMET scores to demonstrate the effectiveness of our method on open sourced models.
@@ 6.1 Meta-Evaluating the Pinpoint Model
Model WMT’22 WMT’23 ASQA Summ Zh-En En-De Zh-En En-De QA Summ PaLM-2 0-shot 66.1 77.0 65.7 75.1 17.6 25.2 Improve 67.7 77.1 67.5 75.9 19.2 25.5 Score-QE 67.5 77.2 67.2 76.3 21.9 26.4 Binary-QE 67.6 77.9 67.3 76.5 21.6 26.0 LLMRefine 68.8 78.6 68.2 76.9 26.1 28.1
@@ 6.4 Iterative Refinement
LLMRefine vs Win Neutral Lose Win lose ratio 0-shot 38% 46% 16% 2.34 Improve 39% 45% 16% 2.44 BLEURT-SCORE-QE 41% 44% 15% 2.79 BLEURT-binary-QE 33% 48% 19% 1.76 score-qe 33% 40% 27% 1.23 binary-qe 34% 48% 18% 1.84
Table 5: We conduct head-to-head human evaluation on LLMRefine against all baselines for single step refinement at WMT22 En-De. We report percentages of win, neutral and lose and win/lose ratio in the table.
@@ 7 Conclusion
In this work, we proposed LLMRefine for incorporating a fine-grained error pinpoint into a text generation pipeline via refinement-feedback model loop. We empirically demonstrated that our fine-grained error pinpoint model with simulated annealing achieves superior performance compared to baseline feedback models and other search algorithms across three text generation tasks. Lastly, our experimental findings are further solidified by the human evaluation study. Humans demonstrate a significant preference for the output of LLMRefine over the baseline outputs.
@@ 8 Limitations
While LLMRefine in theory can be applied to landscape of instruction-fine-tuned large language models, it would be noteworthy to mention that the efficiency may differ when applied with models that lack instruction following capability. Training a large language model with instruction, feedback following ability can be great a future direction to mitigate this issue.
@@ Appendix B Additional Results
Table 14: LLMRefine’s performance on Mistral MoE Jiang et al. (2024) compared to coarse feedback, measured by metricX.
@@
Table 15: LLMRefine’s performance on PALM2 compared to coarse feedback, measured by COMET22.